OpenStudy (anonymous):

solve the equation. Check the solution. -2/x+4=4/1+3

OpenStudy (anonymous):

@campbell_st hey can you help?

OpenStudy (campbell_st):

whats you're solution..?

OpenStudy (anonymous):

i keep getting -8/3 but it isnt right

OpenStudy (campbell_st):

and is the question \[\frac{-2}{x + 4} = \frac{4}{1 + 3}\]

OpenStudy (anonymous):

yes thats correct

OpenStudy (campbell_st):

ok... so that means the right side simplifies to \[\frac{-2}{x +4} = \frac{4}{4}\] which becomes \[\frac{-2}{x +4} = 1\] then multiply both sides by (x +4) which gives \[-2 = x + 4\]

OpenStudy (campbell_st):

does that make sense..?

OpenStudy (anonymous):

yes sorta... but it has to be in fraction form

OpenStudy (campbell_st):

well using the equation you gave me... the answer is x = -6 my concern is that the denominator of the right hand side fraction is perhaps x + 3 and not 1 + 3

OpenStudy (anonymous):

so would it be -11/3?

OpenStudy (campbell_st):

ok... I know the problem

OpenStudy (campbell_st):

the equation should be \[\frac{-2}{x +4} = \frac{4}{x + 3}\] start by cross multiplying to get rid of the denominators \[-2(x + 3) = 4(x + 4)\] when you distribute you get -2x -6 = 4x + 16

OpenStudy (anonymous):

okay then combine like terms which is -6=2x+16

OpenStudy (anonymous):

Correct?

OpenStudy (campbell_st):

add 2x to both sides of the equation -6 = 6x + 16 I think you made a simple error in -2 * 3 = 6

OpenStudy (campbell_st):

now subtract 16 from both sides -22 = 6x

OpenStudy (anonymous):

ohhhhhhh lol

OpenStudy (campbell_st):

now you'll get the correct answer answer...

OpenStudy (anonymous):

thanks!